# DCT-Shield: A Robust Frequency Domain Defense against Malicious Image Editing

### v1

@@ 1 Introduction
Figure 1: A comparison of our proposed method, DCTshield’s, performance compared to SDS(-) [23]. Our method produces highly imperceptible adversarial noise while offering strong protection against edits. Immunized images from DCT-shield are significantly more robust to purification techniques like JPEG-conversion across a wide range of compression qualities.
Current approaches protect images by adding small adversarial noise that disrupts diffusion models, preventing meaningful edits. However, existing methods have key limitations that make them impractical. The added noise is often noticeable when zoomed in, and protection fails when images are saved in low-quality JPEG format. As shown in Fig. 1, prior methods like SDS(-) [23] leave visible artifacts and lack robustness against JPEG compression.
@@
  • We develop a novel immunization algorithm DCT-Shield that adds highly imperceptible noise in the DCT domain, while providing better protection against malicious edits compared to prior works.
  • We demonstrate that incorporating the JPEG pipeline in the noise optimization process allows us to generate quantization-aware perturbations that remain effective even after JPEG compression. Additionally, our algorithm provides users with control over JPEG-aware and JPEG-averse immunization by adjusting quality parameters and thresholds, ensuring adaptable protection against compression-based purification methods.
  • We introduce multiple DCT-Shield variants: the base version offers broad protection against edits, Mask-based DCT-Shield specializes in inpainting defense, and Y-channel DCT-Shield ensures optimal imperceptibility and robustness under high JPEG compression.
@@ 2 Related Work
With advances in diffusion-based image editing models [14, 2, 1] , it has become essential to immunize images against malicious edits. It is worth noting that all prior methods add an adversarial noise in the pixel space of the image so as to minimize some objective, such that the immunized image when processed with an editing model, fails to produce meaningful edits. Typically, the objectives include latent representation of the image (encoder attack) or include the noise predicted by the U-net (diffusion attack) or a combination of both. EditShield [3] performs an encoder attack by maximizing the distance between the latents of the input and immunized images. PhotoGuard [17] performs separate encoder and diffusion attacks by minimizing the distance between the immunized image’s latent and a blank target image’s latent. AdvDM [9] and MIST [8], perform a weighted encoder and diffusion attack with a customized target image. Although they were introduced to prevent copying of art, they are also used to prevent malicious edits. DiffusionGuard [4] performs a diffusion attack to preventing malicious inpainting. Methods like [12, 4, 9, 8], that involve diffusion attacks are typically computationally expensive due to iteratively computing gradients across large U-nets. Diff-Protect [11] employs ideas from score distillation sampling to perform faster and more efficient optimization while performing diffusion attacks. The authors of [11] also provide convincing evidence that VAE encoders are more vulnerable to adversarial attacks compared U-nets in diffusion pipelines.
However, these methods have two major shortcomings. First, the adversarial noise often appears as noticeable patterns in immunized images. Second, they lack guaranteed JPEG robustness, as immunity is typically lost when the image is saved in JPEG format. While increasing the pixel budget can improve JPEG robustness, it comes at the cost of noisy immunized images. Moreover, determining the optimal pixel budget for reliable protection is challenging.
@@ 3.1 Image Editing with Latent Diffusion Models
Latent Diffusion Models (LDMs) [16] have emerged as powerful tools for image editing by leveraging a compressed latent space to perform generative tasks efficiently. Given an image $x_{o}\sim q(x_{0})$ from the real data distribution, an LDM first uses an encoder $\mathcal{E}_{\phi}$ parameterized by $\phi$ to encode $x_{0}$ into a latent variable: $z_{0}=\mathcal{E}_{\phi}(x_{0})$. The forward process follows a Markovian Gaussian noise addition scheme $q(z_{t}|z_{t-1})=\mathcal{N}(z_{t};\sqrt{1-\beta_{t}}z_{t-1},\beta_{t}I)$, where $\beta_{t}$ is a variance schedule controlling the noise magnitude. Over multiple steps, the latent representation $z_{t}$ gradually approaches an isotropic Gaussian Distribution.
To generate or edit an image, the model learns to reverse the noise perturbations in the latent space. The reverse process in parameterized by a neural network $\epsilon_{\theta}(z_{t},t)$ trained to estimate the noise at each step using the following objective function -
$\mathcal{L}(\theta)=\mathbb{E}_{z_{0},t,c}\left\lVert\epsilon_{\theta}(z_{t},t)-\epsilon\right\rVert^{2}$ (1)
@@ 3.2 Adversarial Examples against LDMs
The key idea behind generating adversarial examples is to perturb a clean input $\mathbf{x}$ with some perturbation $\mathbf{\delta}$ such that the perturbation yields unexpected outputs from the LDM. Typically, $\mathbf{\delta}$ is added in the pixel space and an optimization problem is formulated with a specific objective. For example,
@@ 3.3 JPEG Algorithm
The JPEG algorithm [20] compresses images by removing details that are less noticeable to the human eye while maintaining perceptual similarity. We provide a brief description of the JPEG image compression algorithm. The JPEG encoding process consists of the following key steps:
@@
Discrete Cosine Transform (DCT). The core of JPEG compression lies in the application of the Discrete Cosine Transform (DCT). The image is divided into non-overlapping $8\times 8$ pixel blocks, and each block undergoes a transformation from the spatial domain to the frequency domain. The DCT of an $8\times 8$ block $p(x,y)$ is computed as:
$\alpha(u,v)=\frac{1}{4}C(u)C(v)\sum_{x=0}^{7}\sum_{y=0}^{7}p(x,y)g(x,u)g(y,v)$ (3)
@@
Quantization. The DCT coefficients are quantized using a predefined quantization table $Q_{table}$. This step applies lossy compression by reducing the precision of high-frequency components, that are less perceptible to the human eye. The quantized coefficients $Q(u,v)$ are given by
$Q(u,v)=\frac{\alpha(u,v)}{Q_{table}(u,v)*f_{s}.}$ (6)
where, the scaling factor $f_{s}$ is inversely related to the quality factor $Q_{alg}$.
@@
JPEG Decompression The JPEG decoding process reverses the encoding steps to reconstruct an approximation of the original image while maintaining perceptual quality. This involves entropy decoding, followed by de-quantization, where the quantized DCT coefficients are rescaled to approximate the original values. Then, Inverse DCT (IDCT) then converts the frequency-domain data back into spatial-domain pixel values for each 8×8 block. This is followed by color space conversion from YCbCr back into RGB. Finally, the blocks are reassembled to reconstruct the image.
@@ 4 Proposed Method
We propose DCT-Shield in an endeavor to overcome the shortcomings of prior methods (Section. 2) that perform optimization in the pixel space. We are motivated by the JPEG algorithm which essentially modifies image information in the DCT space to produce a JPEG image which is perceptually almost identical to the original image. Since our objective is also to introduce imperceptible modifications in the image, albeit for immunization, we develop DCT-Shield by taking advantage of the JPEG algorithm pipeline. This approach also provides a natural added advantage of in terms of robustness of immunized images against JPEG-purification techniques. We note here, that although we employ the JPEG pipeline for adversarial optimization, the immunized images are free to be saved in any format.
@@ 4.2 Problem Formulation
Given an input image $\mathbf{x}\in\mathbb{R}^{C\times{H}\times{W}}$ we first compute the patch-wise quantized DCT coefficients using the JPEG encode operation:
$\alpha(c)=JPEG_{E}(\mathbf{x}),\quad\alpha(c)\in\mathbb{R}^{n_{p}(c)\times 8\times 8},$
where $c\in\{Y,Cb,Cr\}$ denotes the specific channel and $n_{p}(c)$ represents the number of patches in the channel. $JPEG_{E}$ denotes the sequence of compression operations described in Section 3.3 from RGB-YCbCr conversion to quantizing the DCT-coeffecients.
Our objective is to perturb these quantized coefficients in a way that disrupts the functionality of diffusion-based editing models. To achieve this, we target the VAE encoder within the diffusion pipeline. The key idea is to effectively optimize the perturbation $\delta$ so that the VAE encodes the adversarial image into a distorted or misleading latent representation. Therefore we first obtain an adversarial reconstruction $\mathbf{x^{\prime}}$ by decompressing the perturbed quantized coefficients. Then we compute the latent representation of $\mathbf{x^{\prime}}$ by passing it through a VAE encoder $\mathcal{E}$. Finally, the optimization objective is formulated as minimizing a loss $\mathcal{L}$ as function of $\mathcal{E}(\mathbf{x^{\prime}})$.
@@
Here, $\delta$ denotes the collective noise across all channels, where each channel’s noise $\delta(c)$ is represented as a tensor of dimensions $\mathbb{R}^{n_{p}(c)\times 8\times 8}$. $JPEG_{D}$ denotes the sequence of decompression operations described in Section 3.3. The perturbation $\delta$ in the quantized DCT-coefficients in general is bounded by $\epsilon\in\mathbb{R}$. To ensure JPEG protection, it is crucial to induce at least one quantization level change in the quantized coefficients. Thus, the perturbation must satisfy $|\epsilon|\geq 1$. We use the Projected Gradient Descent (PGD) to solve the optimization problem formulated in Eqn. 7. We also note the fact that we add adversarial perturbations after the quantization step. This is done to allow gradients $\nabla_{\delta}\mathcal{L}$ to back-propagate without having to go through the quantization function, which would otherwise zero out the gradients. Fig. 2 shows a schematic of the optimization process of DCT-Shield.
In our experiments, we employ norm minimization loss on the VAE latents, formulated as $\mathcal{L}(\delta)=\lVert\mathcal{E}(\mathbf{x}^{\prime})\rVert_{2}$. This objective can also be extended to incorporate a target image $\mathbf{x}_{t}$, guiding the VAE latent representation toward a desired target latent. However, based on our empirical observations, norm minimization loss consistently delivers the best performance.
Our proposed framework also allows for a diffusion-based attack, which targets the full diffusion process rather than just the VAE latents. In this approach, the loss function is defined as $\lVert f(\mathbf{x}^{{}^{\prime}})-\mathbf{x}_{t}\rVert_{2}^{2}$, where $f(.)$ denotes the LDM model. However, this method necessitates backpropagation through the diffusion process, making it computationally expensive. Moreover, our experiments indicate that it does not yield significant improvements over the VAE-based attack. This observation aligns with findings from Liu et al. [11], who confirm that VAEs are more susceptible to adversarial attacks than U-Nets within the diffusion process.
Notably, our method reduces the parameter requirement from $O(3HW)$ in pixel-space approaches to $O(3HW/2)$, matching the number of elements in the YCbCr channels after JPEG encoding. Furthermore, some specialized variants of DCT-Shield require only $O(HW)$ parameters while still maintaining strong protection against adversarial attacks.
@@ 4.3 Algorithm Variants
DCT-Shield introduces multiple variants to provide tailored protection against different types of malicious image editing while ensuring high imperceptibility. The base variant applies adversarial noise directly in the DCT domain, optimizing for robust protection across a wide range of editing tasks, including text-guided modifications, inpainting, and outpainting. However, recognizing that different types of editing and attack scenarios require specialized defenses, Mask-based DCT-Shield was developed to focus protection specifically on inpainting tasks. This variant strategically allocates noise to vulnerable regions that are more likely to be edited, enhancing the defense mechanism against localized modifications. Another specialized variant, Y-channel DCT-Shield, is designed to protect against low quality JPEG conversions by restricting adversarial perturbations to the luminance (Y) channel in the YCbCr color space. This ensures that the perturbations remain highly imperceptible, leveraging human visual system properties while maintaining robust protection. Therefore, DCT-Shield’s variants provide a flexible defense, enabling users to balance imperceptibility, robustness, and efficiency based on their needs. Please refer to the detailed algorithm in Section A of the supplementary material.
@@ 5.1 Benchmark Dataset
We use 150 samples from the OmniEdit [22] dataset covering a variety of edit tasks like object addition, removal, replacement and modifications of attributes and environment. For the inpainting task, we constructed a dataset consisting of 56 samples. These samples include images sourced from the PPR10K dataset [10] as well as portrait images collected from the web.
@@ 5.3 Edit Models and Evaluation Metrics
For our experiments, we utilize InstructPix2Pix (IP2P) [2], a widely used diffusion-based editing model, for instruction-based image editing. Additionally, we employ the Stable Diffusion v1.5 Inpainting model [15] for the image inpainting task.
@@
We quantify semantic differences using LPIPS [24], FID [7], PSNR, SSIM [21], and VIFp [19]. For inpainting, we also report CLIPScore [6]. Additionally, we conduct extensive human evaluations for both editing and inpainting. Further details are in the supplementary.

### v2

@@ 1 Introduction
Figure 1: A comparison of our proposed method, DCT-Shield’s, performance compared to SDS(-) [24]. Our method produces highly imperceptible adversarial noise while offering strong protection against edits. Immunized images from DCT-Shield are significantly more robust to purification techniques like JPEG-conversion across a wide range of compression qualities.
Current approaches protect images by adding small adversarial noise that disrupts diffusion models, preventing meaningful edits. However, existing methods have key limitations that make them impractical. The added noise is often noticeable when zoomed in, and protection fails when images are converted to low-quality JPEG format. As shown in Fig. 1, prior methods like SDS(-) [24] leave visible artifacts and lack robustness against JPEG compression.
@@
  • We develop a novel immunization algorithm DCT-Shield that leverages the DCT domain to add highly imperceptible adversarial noise in images, while providing better protection against malicious edits compared to prior works.
  • We demonstrate that incorporating the JPEG pipeline in the noise optimization process allows us to generate quantization-aware perturbations that remain effective even after JPEG-purification. Additionally, DCT-Shield offers adaptable protection, allowing users to tune quality parameters and thresholds to balance the trade-off between purification-robustness and noise-imperceptibility.
  • We introduce multiple DCT-Shield variants: the base version offers broad protection against edits, Mask-based DCT-Shield specializes in inpainting defense, and Y-channel DCT-Shield ensures optimal noise imperceptibility and robustness under high JPEG compression.
@@ 2 Related Work
With advances in diffusion-based image editing models [15, 2, 1] , it has become essential to immunize images against malicious edits. It is worth noting that all prior methods add an adversarial noise in the pixel space of the image so as to minimize some objective, such that the immunized image when processed with an editing model, fails to produce meaningful edits. Typically, the objectives include the image’s latent representation (encoder attack [^1] In this section, ‘attack’ refers to an adversarial attack on the editing model with an immunized image), or the noise predicted by the U-net (diffusion attack) or their combination. EditShield [4] performs an encoder attack by maximizing the distance between the latents of the input and immunized images. PhotoGuard [18] performs separate encoder and diffusion attacks by minimizing the distance between the immunized image’s latent and a blank target image’s latent. AdvDM [10] and MIST [9], perform a weighted encoder and diffusion attack with a customized target image. Although they were introduced to prevent copying of art, they are also used to prevent malicious edits. DiffusionGuard [5] performs a diffusion attack to prevent malicious inpainting. Methods like [13, 5, 10, 9], that involve diffusion attacks are typically computationally expensive due to iteratively computing gradients across large U-nets. Diff-Protect [24] leverages score distillation sampling to enable faster and more efficient optimization during diffusion attacks. The authors of [24] also provide convincing evidence that VAE encoders are more vulnerable to such adversarial attacks compared to U-nets in diffusion pipelines.
However, these methods have two major shortcomings. First, the adversarial noise often appears as noticeable patterns in immunized images. Second, they lack guaranteed JPEG robustness, as immunity is typically lost when the image is compressed with JPEG. While increasing the pixel budget can improve JPEG robustness, it comes at the cost of noisy immunized images. Moreover, determining the optimal pixel budget for reliable protection is challenging.
@@ 3.1 Image Editing with Latent Diffusion Models
Latent Diffusion Models (LDMs) [17] have emerged as powerful tools for image editing by leveraging a compressed latent space to perform generative tasks efficiently. Given an image $x_{o}\sim q(x_{0})$ from the real data distribution, an LDM first uses an encoder $\mathcal{E}_{\phi}$ parameterized by $\phi$ to encode $x_{0}$ into a latent variable: $z_{0}=\mathcal{E}_{\phi}(x_{0})$. The forward process follows a Markovian Gaussian noise addition scheme $q(z_{t}|z_{t-1})=\mathcal{N}(z_{t};\sqrt{1-\beta_{t}}z_{t-1},\beta_{t}I)$, where $\beta_{t}$ is a variance schedule controlling the noise magnitude. Over multiple steps, the latent representation $z_{t}$ gradually approaches an isotropic Gaussian Distribution. To generate an image, the model learns to reverse the noise perturbations in the latent space. The reverse process in parameterized by a neural network $\epsilon_{\theta}(z_{t},t)$ trained to estimate the noise at each step using the following objective function:
$\mathcal{L}(\theta)=\mathbb{E}_{z_{0},t}\left\lVert\epsilon_{\theta}(z_{t},t)-\epsilon\right\rVert^{2}.$ (1)
@@ 3.2 Adversarial Examples against LDMs
The key idea behind generating adversarial examples is to perturb a clean input $\mathbf{x}$ with some perturbation $\mathbf{\delta}$ such that it yields unexpected outputs from the LDM. Typically, $\mathbf{\delta}$ is added in the pixel space and an optimization problem is formulated with a specific objective. For example,
@@ 3.3 JPEG Algorithm
The JPEG algorithm [21] compresses images at a specified quality factor $Q_{alg}$ by removing details that are less noticeable to the human eye while maintaining perceptual similarity. The JPEG encoding process consists of the following key steps:
@@
Discrete Cosine Transform (DCT). The core of JPEG compression lies in the application of the Discrete Cosine Transform (DCT). The image is divided into non-overlapping $\times 8$ 8 pixel blocks, and each block undergoes a transformation from the spatial domain to the frequency domain. The DCT of an $\times 8$ 8 block $p(x,y)$ is computed as:
$\overline{\alpha}(u,v)=\frac{1}{4}C(u)C(v)\sum_{x=0}^{7}\sum_{y=0}^{7}p(x,y)g(x,u)g(y,v)$ (3)
@@
Quantization. The DCT coefficients are quantized using a predefined quantization table $Q_{table}$. This step applies lossy compression by reducing the precision of high-frequency components, that are less perceptible to the human eye. The quantized coefficients $\alpha(u,v)$ are given by
$\alpha(u,v)=\frac{\overline{\alpha}(u,v)}{Q_{table}(u,v)*f_{s}.}$ (6)
where, the scaling factor $f_{s}$ is inversely related to quality factor $Q_{alg}$.
@@
JPEG Decompression The JPEG decoding process reverses the encoding steps to reconstruct an approximation of the original image while maintaining perceptual quality. This involves entropy decoding, followed by de-quantization, where the quantized DCT coefficients are rescaled to approximate the original values. Then, Inverse DCT converts the frequency-domain data back into spatial-domain pixel values for each 8×8 block. This is followed by color space conversion from YCbCr back into RGB. Finally, the blocks are reassembled to reconstruct the image.
@@ 4 Proposed Method
We propose DCT-Shield in an endeavor to overcome the shortcomings of prior methods (Section. 2) that perform optimization in the pixel space. We are motivated by the JPEG algorithm which essentially modifies image information in the DCT space to produce a JPEG image which is perceptually almost identical to the original image. Since our objective is also to introduce imperceptible modifications in the image, albeit for immunization, we develop DCT-Shield by taking advantage of the JPEG algorithm pipeline. This approach also provides a natural added advantage in terms of robustness of immunized images against JPEG-purification techniques. We note here, that although we employ the JPEG pipeline for adversarial optimization, the immunized images are free to be saved in any format.
@@ 4.2 Problem Formulation
Given an input image $\mathbf{x}\in\mathbb{R}^{C\times{H}\times{W}}$ and a JPEG quality setting $Q_{alg}$, we first compute the patch-wise quantized DCT coefficients using the JPEG encode operation as follows:
$\alpha=JPEG_{E}(\mathbf{x};Q_{alg}),\quad\alpha(c)\in\mathbb{Z}^{n_{p}(c)\times 8\times 8},$ (7)
where, $c\in\{Y,Cb,Cr\}$ denotes the specific channel and $n_{p}(c)$ represents the number of patches in the channel. $JPEG_{E}$ denotes the sequence of compression operations described in Section 3.3 from RGB-YCbCr conversion to quantizing the DCT-coefficients.
Our objective is to perturb these quantized coefficients in a way that disrupts the functionality of diffusion-based editing models. To achieve this, we target the VAE encoder within the diffusion pipeline. The key idea is to effectively optimize the perturbation $\delta$ so that the VAE encodes the adversarial image into a distorted or misleading latent representation. Therefore we first obtain an adversarial reconstruction $\mathbf{x^{\prime}}$ by decompressing the perturbed quantized coefficients. Then we compute the latent representation of $\mathbf{x^{\prime}}$ by passing it through a VAE encoder $\mathcal{E}$. Finally, the optimization objective is formulated as minimizing a loss $\mathcal{L}$ as a function of $\mathcal{E}(\mathbf{x^{\prime}})$, i.e.,
@@
Here, $\delta$ denotes the collective noise across all channels, where each channel’s noise $\delta(c)$ is represented as a tensor of dimensions $\mathbb{R}^{n_{p}(c)\times 8\times 8}$. $JPEG_{D}$ denotes the sequence of decompression operations described in Section 3.3. The perturbation $\delta$ in the quantized DCT-coefficients in general is bounded by $\epsilon\in\mathbb{R}^{+}$. To ensure JPEG protection, it is crucial to induce at least one quantization level change in the quantized coefficients. Thus, the perturbation must satisfy $\epsilon\geq 1$. We use the Projected Gradient Descent (PGD) to solve the optimization problem formulated in Eqn. 8. We also note the fact that we add adversarial perturbations after the quantization step. This is done to allow gradients $\nabla_{\delta}\mathcal{L}$ to back-propagate without having to go through the quantization function, which would otherwise zero out the gradients. Fig. 2 shows a schematic of the optimization process of DCT-Shield.
In our experiments, we employ a norm minimization loss on the VAE latents, formulated as $\mathcal{L}(\delta)=\lVert\mathcal{E}(\mathbf{x}^{\prime})\rVert_{2}$. This objective can also be extended to incorporate a target image $\mathbf{x}_{t}$, guiding the VAE latent representation toward a desired target latent. However, based on our empirical observations, norm minimization loss consistently delivers the best performance.
Our proposed framework also allows for a diffusion-based objective, which targets the full diffusion process rather than just the VAE latents. In this approach, the loss function is defined as $\lVert f(\mathbf{x}^{{}^{\prime}})-\mathbf{x}_{t}\rVert_{2}^{2}$, where $f(.)$ denotes the LDM model. However, this method necessitates backpropagation through the diffusion process, making it computationally expensive. Moreover, our experiments indicate that it does not yield significant improvements over the VAE-based objective. This observation aligns with findings from Liu et al. [24], who confirm that VAEs are more susceptible to adversarial attacks than U-Nets within the diffusion process.
Notably, our method reduces the parameter requirement from $O(3HW)$ in pixel-space approaches to $O(3HW/2)$, matching the number of elements in the YCbCr channels after JPEG encoding. Furthermore, some specialized variants of DCT-Shield require only $O(HW)$ parameters while still maintaining strong protection against malicious edits.
@@ 4.3 Algorithm Variants
DCT-Shield introduces multiple variants to provide tailored protection against different types of malicious image editing while ensuring high imperceptibility. The base variant applies adversarial noise directly in the DCT domain, optimizing for robust protection across a wide range of editing tasks, including text-guided modifications, inpainting, and outpainting. However, recognizing that different types of editing and attack scenarios require specialized defenses, Mask-based DCT-Shield was specifically tailored for inpainting tasks. This variant strategically allocates noise to vulnerable regions that are more likely to be edited, enhancing the defense mechanism against localized modifications. Another specialized variant, Y-channel DCT-Shield, is designed to protect against low quality JPEG conversions by restricting adversarial perturbations to the luminance (Y) channel in the YCbCr color space. This ensures that the perturbations remain highly imperceptible while maintaining robust protection. Therefore, DCT-Shield’s variants provide a flexible defense, enabling users to balance noise-imperceptibility, protection-robustness, and efficiency based on their needs. Please refer to the detailed algorithm in Section A of the supplementary material.
@@ 5.1 Benchmark Dataset
We use 150 samples from the OmniEdit [23] dataset covering a variety of edit tasks like object addition, removal, replacement and modifications of attributes and environment. For the inpainting task, we constructed a dataset consisting of 56 samples. These samples include images sourced from the PPR10K dataset [11] as well as portrait images collected from the web. Further details and dataset samples are shown in Section B of the supplementary material.
@@ 5.3 Edit Models and Evaluation Metrics
For our experiments, we utilize InstructPix2Pix (IP2P) [2], a widely used diffusion-based editing model, for instruction-based image editing. For the inpainting task, we employ the Stable Diffusion Inpainting 1.0 model [16]. Additionally, we demonstrate the cross-model transferability of our approach in Section D.5 of the supplementary material.
@@
We quantify semantic differences using LPIPS [25], FID [8], PSNR, SSIM [22], and VIFp [20]. For inpainting, we also report CLIPScore [7]. Additionally, we conduct extensive human evaluations for both editing and inpainting. Further details are in the Section C of the supplementary.